Each front tire on a particular type of vehicle is supposed to be filled to a pressure of 26 psi. Suppose the actual air presPlease solve a and b
Answer #1
a) E(X)= ∫22322232 kx(x2+y2) dx dy = 27.3052 =E(Y)
E(X2)= ∫22322232 kx2(x2+y2) dx dy = 753.8493
Var(X)=E(X2)-(E(X))2 = 8.2779 =VAR(Y)
E(XY)= ∫22322232 kxy(x2+y2) dx dy = 745.4783
a) therefore Cov(X,Y)=E(XY)-E(X)*E(Y)= -0.0931
b) and correlation coefficient ρ =Cov(X,Y)/(σx*σy)= -0.0112